OpenStudy (anonymous):

Express the 4th-Order differential equation as a system of first order differential equations using matrix notation

OpenStudy (anonymous):

\[y^{(4)}+\frac{ y^{(3)} }{ t^{2} }+\sin(ty'')+e^ty=t^2\]

OpenStudy (anonymous):

i know to do this, we need to set the derivative of y to another variable, but i'm confused on the steps to do so, and how that will actually help solve this too

OpenStudy (anonymous):

do i just set something like: \[y^{(3)} = x_1 \] and\[y^{(4)} = x_1'\]